OpenStudy (anonymous):

In the formula below, represents the electric power of a light bulb measured in watts, is the voltage, and is the current in amperes. A 150-watt light bulb has a current of 6 amperes and a voltage of 25 volts. If the voltage of the 150-watt light bulb doubles, what will happen to the current? The current will stay the same. The current will double. The current will decrease by one-fourth. The current will decrease by half.

OpenStudy (luckycoins888):

The current will stay the same.

OpenStudy (radar):

I will double, E=IR. For all practical purposes the resistance of the bulb is a constant. The resistance is, using R=E/I = 25/6 = 4 1/6 Ohm Applying a voltage of 50 volts (doubling the voltage) Compute the current using I= E/R = 50/4.166666 = 11.999999 or doubles also from the former value of 6 amperes.